# Velocity effects slightly mitigating the quantumness degradation of an Unruh-DeWitt detector

### v1

@@ II.2.1 Non-relativistic velocity regime
It is important to note that when we take $w=0$ in Eqs. (18) and (23), we obtain the standard case without velocity effects sriramkumar1996finite ; Pedro2024robustness . The expressions Eq. (18) and Eq. (23) present slight modifications of the thermal spectrum, moreover, the dependence of $w$ highlights the velocity effect on the detector transition probability rates.
II.2.2 Ultra-relativistic velocity regime
Now, let us examine the case in which the detector follows the trajectory given in Eq. (10), in the ultra-relativistic regime, namely in the limit $w\to\infty$. In this context, the Wightman function of the massless scalar field, as expressed in Eq. (12), which is suppressed as
$\mathcal{W}^{+}_{w}(\Delta\tau)=-\frac{a^{2}}{16\pi^{2}}\left[\sinh^{2}{\left(\frac{a\Delta\tau}{2}-i\epsilon a\right)}\right]^{-1}\frac{1}{w^{4}},$ (27)
@@ II.2.2 Ultra-relativistic velocity regime
$\overline{\mathcal{R}}^{-}_{w}(\infty)$ $=$ $\frac{1}{2\pi}\frac{1}{(\mbox{\rm e}^{2\pi/\overline{a}}-1)}\frac{1}{w^{4}},$ (28)
@@
In this way, it is clarified through Eqs. (28) and (29), that the Unruh effect is suppressed when the four-component of the velocity reaches the ultra-relativistic regime ($w\to\infty)$. It is important to emphasize that since the Unruh effect is suppressed in the ultra-relativistic regime, then the UDW detector will not respond, and consequently, we will not have the influence of acceleration radiation on accelerated quantum systems.
III Velocity effects on an accelerated single-qubit
III.1 The theoretical model
@@
where $\alpha$ and $\beta$ are complex amplitudes given by $\alpha=\mbox{\rm e}^{i\frac{\varphi}{2}}\cos{\frac{\theta}{2}}$, $\beta=\mbox{\rm e}^{-i\frac{\varphi}{2}}\sin{\frac{\theta}{2}}$, and they satisfy the relationship $|\alpha|^{2}+|\beta|^{2}=1$. See that here, $\theta\in[0,\pi]$ and $\varphi\in[0,2\pi]$ are the polar and azimuthal angles of the Bloch sphere jazaeri2019review ; kasirajan2021quantum . In this way, the density matrix $\hat{\rho}^{\mathrm{in}}_{D}=|\psi_{D}\rangle\langle\psi_{D}|$, using the Eq. (30), read as
@@
$\hat{\rho}^{\mathrm{out}}$ $=$ $\mathcal{\hat{U}}^{(0)}\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(0)^{\dagger}}+\mathcal{\hat{U}}^{(1)}\hat{\rho}_{\mathrm{in}}+\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(1)^{\dagger}}+\mathcal{\hat{U}}^{(1)}\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(1)^{\dagger}}+\mathcal{\hat{U}}^{(2)}\hat{\rho}_{\mathrm{in}}+\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(2)^{\dagger}}+\mathcal{O}(\lambda^{3}).$ (32)
@@
$\mathcal{\hat{U}}^{(0)}$ $=$ $\mathbb{I};$ (34)
$\mathcal{\hat{U}}^{(1)}$ $=$ $-i\lambda\int_{-\infty}^{\infty}d\tau\chi(\tau)\mu(\tau)\phi[x(\tau)];$ (35)
@@
$\hat{\rho}_{D}$ $=$ $\left\{\cos^{2}{\frac{\theta}{2}}+\lambda^{2}\left[\sin^{2}{\frac{\theta}{2}}\mathcal{F}^{-}-2\cos^{2}{\frac{\theta}{2}}\mathbf{Re}\left(\mathcal{G}^{-}\right)\right]\right\}|g\rangle\langle g|$ (41)
$+$ $\left\{\sin^{2}{\frac{\theta}{2}}+\lambda^{2}\left[\cos^{2}{\frac{\theta}{2}}\mathcal{F}^{+}-2\sin^{2}{\frac{\theta}{2}}\mathbf{Re}\left(\mathcal{G}^{+}\right)\right]\right\}|e\rangle\langle e|$
@@
It should be emphasized that the reduced density matrix given in Eq. (41) must possess unit trace. In order to satisfy this requirement, the terms proportional to $\lambda^{2}$ cancel as a direct consequence of the normalization condition. Accordingly, one obtains
$\mathbf{Re}(\mathcal{G}^{-})=\frac{1}{2}\left(\mathcal{F}^{-}\sin^{2}{\frac{\theta}{2}}+\mathcal{F}^{+}\cos^{2}{\frac{\theta}{2}}\right).$ (42)
This is a consequential relation of the trace property of the density matrix.
III.2 Quantum coherence
Quantum coherence is a fundamental property of quantum systems, manifested through the existence of superposition states that enable interference between distinct eigenstates Leggett1980 . More precisely, it is characterized by the preservation of relative phases among the components of a superposed quantum state, thereby facilitating essential quantum phenomena such as interference and entanglement streltsov2015measuring .
Within this context, quantum optical methods represent a crucial set of techniques for the manipulation and control of coherence Glauber1963coherent ; sudarshan1963 . In the particular case of a two-level quantum system, the coherence between the states $|g\rangle$ and $|e\rangle$ can be quantified by means of the $l^{1}$ norm quantum coherence, defined as the sum of the absolute values of the off-diagonal elements of the system’s density matrix Baumgratz2014QuantifyingCoherence , namely:
@@ III.2 Quantum coherence
$\mathcal{Q}^{l^{1}}=|\sin{\theta}|\left\{1-\lambda\left[2\mathbf{Re}(\mathcal{G}^{-})-\mathcal{C}^{-}\cos{(2\varphi)}\right]\right\}+\mathcal{O}(\lambda^{4}).$ (44)
Using the relation described in Eq. (42), knowing that $\mathcal{F}^{\pm}=\sigma\overline{\mathcal{R}}^{\pm}$, and considering a long interaction time [^1] This regime is necessary to ensure that the detector has enough time to interact with the quantum field., namely $\sigma\ll 1$, we have
$\mathcal{Q}^{l^{1}}_{w}=|\sin{\theta}|\left\{1-\left[\frac{1}{2\pi}-w^{2}F(\overline{a})(\mbox{\rm e}^{2\pi/\overline{a}}-1)\right]\frac{\sigma\lambda^{2}}{\mbox{\rm e}^{2\pi/\overline{a}}-1}\left(\sin^{2}{\frac{\theta}{2}}+\mbox{\rm e}^{2\pi/\overline{a}}\cos^{2}{\frac{\theta}{2}}\right)\right\},$ (45)
@@
$\mathcal{Q}^{l^{1}}_{0}=|\sin{\theta}|\left\{1-\frac{1}{2\pi}\frac{\sigma\lambda^{2}}{\mbox{\rm e}^{2\pi/\overline{a}}-1}\left(\sin^{2}{\frac{\theta}{2}}+\mbox{\rm e}^{2\pi/\overline{a}}\cos^{2}{\frac{\theta}{2}}\right)\right\},$ (46)
for more details on the expression of Eq. (46) see refs. barros2024detecting ; pedro2025mitigating taking the corresponding limits for this.
III.3 Numerical results
@@
IV Velocity effects on the duality relation
IV.1 Quantum interferometric circuit
In this section, we investigate the quantum scattering circuit. The circuit is structured as follows: initially, a single qubit is prepared in a well-defined initial state. Subsequently, the application of the first Hadamard gate transforms the qubit into a coherent superposition. Thereafter, a phase accumulation takes place through the action of the phase-shift gate $\hat{\alpha}$. Following this step, the unitary operator $\hat{\mathcal{U}}$ [Eq. (33)] induces a linear interaction, in a controlled manner, between the detector and the quantum field. Finally, the qubit is subjected to a second Hadamard operation, after which projective measurements on the system are performed.
@@ IV.1 Quantum interferometric circuit
$\hat{\rho}_{I}$ $=$ $\Big\{\cos^{2}{\frac{\alpha}{2}}+\frac{\lambda^{2}}{4}\Big[\mathcal{F}^{-}+\mathcal{F}^{+}-8\mathbf{Re}(\mathcal{G}^{-})\cos^{2}{\frac{\alpha}{2}}+2\mathcal{C}^{-}\cos{\alpha}\Big]\Big\}|g\rangle\langle g|+$ (49)
$+$ $\Big\{\sin^{2}{\frac{\alpha}{2}}+\frac{\lambda^{2}}{4}\Big[\mathcal{F}^{-}+\mathcal{F}^{+}-8\mathbf{Re}(\mathcal{G}^{-})\sin^{2}{\frac{\alpha}{2}}-2\mathcal{C}^{-}\cos{\alpha}\Big]\Big\}|e\rangle\langle e|+$
@@
$+$ $\Big\{+\frac{i}{2}\sin{\alpha}+\frac{\lambda^{2}}{4}\Big[\mathcal{F}^{-}-\mathcal{F}^{+}-2i\sin{\alpha\left(2\mathbf{Re}(\mathcal{G}^{-})-\mathcal{C}^{-}\right)}\Big]\Big\}|g\rangle\langle e|,$
where it must obey the following relation,
$\mathbf{Re}(\mathcal{G}^{-})=\frac{1}{4}\left(\mathcal{F}^{-}+\mathcal{F}^{+}\right),$ (50)
for the terms proportional to $\lambda^{2}$ in the unitary matrix to vanish due to the normalization condition.
IV.2 Visibility and quantum coherence
@@
IV.3 Which-path distinguishability circuit
Now, we consider a modified configuration aimed at extracting which-path information. Specifically, the second Hadamard gate is removed from the quantum interferometric circuit, and two detectors are incorporated. The introduction of the path detectors allows the interaction between the qubit and a massless scalar field to reveal the trajectory followed by the qubit. Specifically, the internal state of the qubit is initialized in the state $|g\rangle$ for qubits registered along path A by detector A. Conversely, qubits that are not registered by detector A are inferred to have propagated along path B. In this manner, the which-path information is effectively encoded in the internal states of the qubit.
@@ IV.3 Which-path distinguishability circuit
IV.4 Complementarity relation
The wave–particle duality wootters1979complementarity ; greenberger1988simultaneous encapsulates the intrinsic trade-off between these two physical attributes. In particular, any attempt to obtain which-path information (i.e., an increase in the distinguishability $\mathcal{D}_{w}$) necessarily leads to a reduction in the coherence of the interference pattern (i.e., a decrease in the visibility $\mathcal{V}_{I,w}$), and conversely bohr1928quantum ; busch1985indeterminacy . This fundamental trade-off is rigorously formalized by the inequality originally derived in the seminal works of Englert and Zeilinger Englert1996 ; Englert1999 ; Zeilinger1999Experiment , expressed as

### v2

@@ 2.2.1 Non-relativistic velocity regime
It is important to note that when we take $w=0$ in Eqs. (18) and (23), we obtain the standard case without velocity effects Sriramkumar and Padmanabhan (1996); Barros et al. (2024). The expressions Eq. (18) and Eq. (23) present slight modifications of the thermal spectrum, moreover, the dependence of $w$ highlights the velocity effect on the detector transition probability rates.
2.2.2 Ultra-relativistic velocity regime
Now, let us examine the case in which the detector follows the trajectory given in Eq. (10), in the ultra-relativistic regime, namely in the limit $w\gg 1$. In this context, we derive in detail in Appendix A the expansion for $w\ll 1$ of the Wightman function, as expressed in Eq. (12), which is written as
$\mathcal{W}_{w}^{+}(\Delta\tau)$ $\approx$ $-\frac{a^{2}}{16\pi^{2}}\frac{1}{w^{4}}\Bigg[\sinh^{2}{\left(\frac{a\Delta\tau}{2w}-\frac{i\epsilon a}{w^{2}}\right)}\Bigg]^{-1}+\mathcal{O}(w^{-6}),$ (27)
@@ 2.2.2 Ultra-relativistic velocity regime
$\overline{\mathcal{R}}^{-}_{w}(\infty)$ $=$ $\frac{1}{2\pi}\frac{1}{(\mbox{\rm e}^{2\pi w/\overline{a}}-1)}\frac{1}{w^{2}},$ (28)
@@
In Fig 1, we plot the excitation probability rate $\overline{\mathcal{R}}^{-}_{w}(\infty)$ as a function of different parameters, in order to analyze the effect of the four-velocity component on the detector response behavior. Note that in both Fig. 1(a) and Fig. 1(b), as the parameter $w$ increases, the detector excitation rate decreases rapidly until it becomes zero. It is important to emphasize that for this regime, the detector response is suppressed even when considering small values for $w$ [see Fig. 1(a)]. Furthermore, see in Fig. 1(b) that for $w\geq 4.00$ the detector does not respond, and thus it is clear that using values of the type $w\gg 1$ the detector will not respond.
In this way, it is clarified through Eqs. (28) and (29), that the Unruh effect is suppressed when the four-component of the velocity reaches the ultra-relativistic regime ($w\gg 1$). It is important to emphasize that since the Unruh effect is suppressed in the ultra-relativistic regime, then the UDW detector will not respond, and consequently, we will not have the influence of acceleration radiation on accelerated quantum systems.
Figure 1: Excitation probability rate $\overline{\mathcal{R}}^{-}_{w}(\infty)$ for the regime where $w\gg 1$: (a) as a function of the parameter $w$ for different values of the parameter $\overline{a}$ and (b) as a function of the parameter $\overline{a}$ for different values of the parameter $w$.
3 Velocity effects on an accelerated single-qubit
3.1 The theoretical model
@@
where $\alpha$ and $\beta$ are complex amplitudes given by $\alpha=\mbox{\rm e}^{i\frac{\varphi}{2}}\cos{\frac{\theta}{2}}$, $\beta=\mbox{\rm e}^{-i\frac{\varphi}{2}}\sin{\frac{\theta}{2}}$, and they satisfy the relationship $|\alpha|^{2}+|\beta|^{2}=1$. See that here, $\theta\in[0,\pi]$ and $\varphi\in[0,2\pi]$ are the polar and azimuthal angles of the Bloch sphere Jazaeri et al. (2019); Kasirajan (2021). In this way, the density matrix $\hat{\rho}^{\mathrm{in}}_{D}=|\psi_{D}\rangle\langle\psi_{D}|$, using the Eq. (30), read as
@@
$\hat{\rho}^{\mathrm{out}}$ $=$ $\mathcal{\hat{U}}^{(0)}\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(0)^{\dagger}}+\mathcal{\hat{U}}^{(1)}\hat{\rho}_{\mathrm{in}}+\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(1)^{\dagger}}+\mathcal{\hat{U}}^{(1)}\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(1)^{\dagger}}+\mathcal{\hat{U}}^{(2)}\hat{\rho}_{\mathrm{in}}+\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(2)^{\dagger}}+\mathcal{O}(\lambda^{3}).$
@@
$\mathcal{\hat{U}}^{(0)}$ $=$ $\mathbb{I},$ (34)
$\mathcal{\hat{U}}^{(1)}$ $=$ $-i\lambda\int_{-\infty}^{\infty}d\tau\chi(\tau)\mu(\tau)\phi[x(\tau)],$ (35)
@@
$\hat{\rho}_{D}$ $=$ $\left[\cos^{2}{\frac{\theta}{2}}+\lambda^{2}\left(\mathcal{F}^{+}\sin^{4}{\frac{\theta}{2}}-\mathcal{F}^{-}\cos^{4}{\frac{\theta}{2}}\right)\right]|g\rangle\langle g|$ (41)
$+$ $\left[\sin^{2}{\frac{\theta}{2}}+\lambda^{2}\left(\mathcal{F}^{-}\cos^{4}{\frac{\theta}{2}}-\mathcal{F}^{+}\sin^{4}{\frac{\theta}{2}}\right)\right]|e\rangle\langle e|$
@@
It should be emphasized that the reduced density matrix given in Eq. (41) must possess unit trace. In order to satisfy this requirement, the terms proportional to $\lambda^{2}$ cancel as a direct consequence of the normalization condition. This is a consequential relation of the trace property of the density matrix.
3.2 Quantum coherence
Quantum coherence is a fundamental property of quantum systems, manifested through the existence of superposition states that enable interference between distinct eigenstates Leggett (1980). More precisely, it is characterized by the preservation of relative phases among the components of a superposed quantum state, thereby facilitating essential quantum phenomena such as interference and entanglement Streltsov et al. (2015).
Within this context, quantum optical methods represent a crucial set of techniques for the manipulation and control of coherence Glauber (1963); Sudarshan (1963). In the particular case of a two-level quantum system, the coherence between the states $|g\rangle$ and $|e\rangle$ can be quantified by means of the $l^{1}$ norm quantum coherence, defined as the sum of the absolute values of the off-diagonal elements of the system’s density matrix Baumgratz et al. (2014), namely:
@@ 3.2 Quantum coherence
$\mathcal{Q}^{l^{1}}=|\sin{\theta}|\left\{1-\lambda^{2}\left[\mathcal{F}^{+}\sin^{2}{\frac{\theta}{2}}+\mathcal{F}^{-}\cos^{2}{\frac{\theta}{2}}-\mathcal{C}^{-}\cos{(2\varphi)}\right]\right\}+\mathcal{O}(\lambda^{4}).$ (43)
Using the relation $\mathcal{F}^{\pm}=\sigma\overline{\mathcal{R}}^{\pm}$, and considering a long interaction time [^1] This regime is necessary to ensure that the detector has enough time to interact with the quantum field., namely $\sigma\gg 1$, we have
$\mathcal{Q}^{l^{1}}_{w}=|\sin{\theta}|\left\{1-\left[\frac{1}{2\pi}-w^{2}F(\overline{a})(\mbox{\rm e}^{2\pi/\overline{a}}-1)\right]\frac{\sigma\lambda^{2}}{\mbox{\rm e}^{2\pi/\overline{a}}-1}\left(\cos^{2}{\frac{\theta}{2}}+\mbox{\rm e}^{2\pi/\overline{a}}\sin^{2}{\frac{\theta}{2}}\right)\right\},$ (44)
@@
$\mathcal{Q}^{l^{1}}_{0}=|\sin{\theta}|\left\{1-\frac{1}{2\pi}\frac{\sigma\lambda^{2}}{\mbox{\rm e}^{2\pi/\overline{a}}-1}\left(\cos^{2}{\frac{\theta}{2}}+\mbox{\rm e}^{2\pi/\overline{a}}\sin^{2}{\frac{\theta}{2}}\right)\right\},$ (45)
for more details on the expression of Eq. (45) see refs. Barros and Costa (2024a); Barros et al. (2025b) taking the corresponding limits for this.
3.3 Numerical results
@@
4 Velocity effects on the duality relation
Wave–particle duality constitutes a foundational principle of quantum theory and is quantitatively expressed through complementarity relations that connect interference visibility with path distinguishability Englert (1996); Wootters and Zurek (1979); Greenberger and Yasin (1988); Bohr and others (1928); Busch (1985); Englert (1999); Zeilinger (1999). In interferometric settings, the visibility provides a measure of the coherence between alternative quantum paths, whereas the distinguishability quantifies the degree of which-path information that may, in principle, be obtained by correlating the interfering system with an external degree of freedom Jaeger et al. (1995); Englert (1996); Miniatura et al. (2007); Kolář et al. (2007); Bera et al. (2015). These quantities satisfy well-established duality relations, which furnish an operational framework for analyzing the manner in which quantum coherence is diminished due to the transfer of information to the environment.
Within the framework of relativistic quantum systems, and particularly in the case of UDW detectors, this approach provides a natural generalization of the single-detector analysis developed in the preceding sections. Although the response of an individual detector already contains nontrivial information concerning motion-induced effects and field correlations Unruh (1981); Fulling (1973); Davies (1975); Letaw (1981); Abdolrahimi (2014), the incorporation of the detector’s internal degree of freedom into a quantum interferometric circuit enables a direct investigation of how relativistic motion influences the interplay between coherence and information extraction Barros et al. (2024, 2025a, 2025b); Costa et al. (2020); Gooding et al. (2020); Lopes et al. (2021). From this viewpoint, the interferometric formulation should not be regarded as an independent problem, but rather as an operationally transparent framework in which velocity-dependent modifications of the detector–field interaction are manifested through quantitative variations in interference visibility and path distinguishability, thereby naturally embedding the analysis of duality relations into the present study.
4.1 Quantum interferometric circuit
One indirect strategy for probing the properties of a physical system consists in the implementation of a quantum scattering circuit, which exploits quantum interferometry Ramsey (1950) to extract information about a specific subsystem by analyzing the measurable properties of an auxiliary probe. Quantum scattering circuits have found a wide range of applications, including tests of the Leggett–Garg inequality Souza et al. (2011), the measurement of correlation functions in simulations of the Fano–Anderson model Negrevergne et al. (2005), the determination of discrete Wigner functions Leonhardt (1995); Miquel et al. (2002), and the experimental reconstruction of work distributions Batalhão et al. (2014), among others. Such a circuit is characterized by a controlled interaction between a single qubit, which acts as a probe, and the system under investigation.
In this section, we investigate the quantum scattering circuit (see a schematic representation in Fig.  5). The circuit is structured as follows: initially, a single qubit is prepared in a well-defined initial state. Subsequently, the application of the first Hadamard gate transforms the qubit into a coherent superposition. Thereafter, a phase accumulation takes place through the action of the phase-shift gate $\hat{\alpha}$. Following this step, the unitary operator $\hat{\mathcal{U}}$ [Eq. (33)] induces a linear interaction, in a controlled manner, between the detector and the quantum field. Finally, the qubit is subjected to a second Hadamard operation, after which projective measurements on the system are performed.
Figure 5: Schematic representation of the quantum interferometric circuit.
@@ 4.1 Quantum interferometric circuit
$\hat{\rho}_{I}$ $=$ $\Big[\cos^{2}{\frac{\alpha}{2}}-\frac{\lambda^{2}}{4}\Big(\mathcal{F}^{-}+\mathcal{F}^{+}-2\mathcal{C}^{-}\Big)\cos{\alpha}\Big]|g\rangle\langle g|+$ (48)
$+$ $\Big[\sin^{2}{\frac{\alpha}{2}}+\frac{\lambda^{2}}{4}\Big(\mathcal{F}^{-}+\mathcal{F}^{+}-2\mathcal{C}^{-}\Big)\cos{\alpha}\Big]|e\rangle\langle e|+$
@@
$+$ $\Big\{+\frac{i}{2}\sin{\alpha}+\frac{\lambda^{2}}{4}\Big[\mathcal{F}^{-}-\mathcal{F}^{+}-2i\sin{\alpha\left(2\mathbf{Re}(\mathcal{G}^{-})-\mathcal{C}^{-}\right)}\Big]\Big\}|g\rangle\langle e|.$
Note that the unitarity condition is satisfied. It is easy to see that the diagonal terms proportional to $\lambda^{2}$ vanish when we compute the trace of this matrix.
4.2 Visibility and quantum coherence
@@
4.3 Which-path distinguishability circuit
Now, we consider a modified configuration aimed at extracting which-path information. Specifically, the second Hadamard gate is removed from the quantum interferometric circuit, and two detectors are incorporated (as shown in Fig.  6). The introduction of the path detectors allows the interaction between the qubit and a massless scalar field to reveal the trajectory followed by the qubit. Specifically, the internal state of the qubit is initialized in the state $|g\rangle$ for qubits registered along path A by detector A. Conversely, qubits that are not registered by detector A are inferred to have propagated along path B. In this manner, the which-path information is effectively encoded in the internal states of the qubit.
Figure 6: Schematic representation of the which-path distinguishability circuit.
@@ 4.3 Which-path distinguishability circuit
4.4 Complementarity relation
The wave–particle duality Wootters and Zurek (1979); Greenberger and Yasin (1988) encapsulates the intrinsic trade-off between these two physical attributes. In particular, any attempt to obtain which-path information (i.e., an increase in the distinguishability $\mathcal{D}_{w}$) necessarily leads to a reduction in the coherence of the interference pattern (i.e., a decrease in the visibility $\mathcal{V}_{I,w}$), and conversely Bohr and others (1928); Busch (1985). This fundamental trade-off is rigorously formalized by the inequality originally derived in the seminal works of Englert and Zeilinger Englert (1996, 1999); Zeilinger (1999), expressed as